# Quantifying Privacy Leakage in Split Inference via Fisher-Approximated Shannon Information Analysis

### v1

@@ Appendix D Examining Defense mechanisms
Tan et al. (Tan et al., 2024) have modeled the estimation error of the DRA adversaries who aim at reconstructing the training dataset $D\in\mathbb{R}^{d}$ of one client from the model parameters $W\in\mathbb{R}^{m}$ in the federated learning scenarios by giving Theorem 1 in (Tan et al., 2024):
Theorem E.1.
(Theorem 1 in (Tan et al., 2024)). For any random variable $D\in\mathbb{R}^{d}$ and $W\in\mathbb{R}^{m}$, we have
@@
By simply applying Theorem E.1 in the privacy quantification in smashed data about raw input in split inference settings, we reach Theorem 4.2. For better understanding, we follow (Tan et al., 2024) to give the proof details of Theorem 4.2 as follows:
@@
Lemma E.2.
(Lemma 2 in (Tan et al., 2024)). For any $d_{x}$-dimensional semi-positive definite matrix $A$ ($A\in\mathbb{R}^{d_{x}\times d_{x}}$), we have:
@@
Let $X$ be the raw input and $\hat{X}$ be the reconstructed input, we have:
@@
$=$ $\mathbb{E}_{x\sim p(X)}H(\hat{X}|X=x)$ (32)
$\leq$ $\mathbb{E}_{x\sim p(X)}\left[\frac{d_{x}}{2}\log(2\pi e)+\frac{1}{2}\log\det(Cov(\hat{X}|X=x))\right]$ (33)
$=$ $\frac{d_{x}}{2}\log(2\pi e)+\mathbb{E}_{x\sim p(X)}\left[\frac{1}{2}\log\det(Cov(\hat{X}|X=x))\right]$ (34)
$\leq$ $\frac{d_{x}}{2}\log(2\pi e)+\mathbb{E}_{x\sim p(X)}\left[\frac{d_{x}}{2}\log\frac{tr(Cov(\hat{X}|X=x))}{d_{x}}\right]$ (35)
$=$ $\frac{d_{x}}{2}\log(2\pi e)+\mathbb{E}_{x\sim p(X)}\left[\frac{d_{x}}{2}\log\frac{\mathbb{E}_{p(X,\hat{X})}\left[||\hat{X}-X||^{2}|X=x\right]}{d_{x}}\right]$ (36)
$\leq$ $\frac{d_{x}}{2}\log(2\pi e)+\frac{d_{x}}{2}\log\frac{\mathbb{E}_{x\sim p(X)}\left[\mathbb{E}_{p(X,\hat{X})}\left[||\hat{X}-X||^{2}|X=x\right]\right]}{d_{x}}$ (37)
$=$ $\frac{d_{x}}{2}\log(2\pi e)+\frac{d_{x}}{2}\log\frac{\mathbb{E}_{p(X,\hat{X})}\left[||\hat{X}-X||^{2}\right]}{d_{x}}.$ (38)
The first inequality holds because the Gaussian distribution is the maximum entropy distribution among distributions with the same mean value and covariance. The second inequality depends on Lemma E.2. The Equation 36 holds because,
@@
Lemma F.1.
(Theorem 9 in (Scarlett & Cevher, 2019)). Under the minimax estimation setup in Section 4.3, fix $\epsilon>0$, and index $V\in\mathcal{V}$ is drawn from a prior distribution $P_{V}$. $X_{V}=\{x_{v_{1}},\dots,x_{v_{|\mathcal{V}|}}\}$ drawn from a probability distribution $P_{X|V}$ be a finite subset of $\mathcal{X}$ such that
@@
In Lemma F.1, the $X_{V}$ is identified by $V$, so we let both $V$ and $X_{V}$ be the raw input $X$ in split inference. In this case, given the $V=X$, we have $X_{V}=V=X$, where the $X_{V}$ is uniquely identified by $V$. Then, we let $Z$ be the smashed data $Z$ and both $\hat{X}$ and $\hat{V}$ be the reconstructed input $\hat{X}$ in split inference system, and we have the Markov chain
@@
Follow (Noorbakhsh et al., 2024), we assume $X$ is uniform over $\mathcal{X}$. Then by applying Lemma F.1 and using $I(X;\hat{X})=H(\hat{X})-H(\hat{X}|X)$, we reach out Theorem 4.4. ∎

### v2

@@ 2.2 For big models (DistilBert), we want to save the memory cost:
Tan et al. [sec24-defending-FL-ita] have modeled the estimation error of the DRA adversaries who aim at reconstructing the training dataset $D\in\mathbb{R}^{d}$ of one client from the model parameters $W\in\mathbb{R}^{m}$ in the federated learning scenarios by giving Theorem 1 in [sec24-defending-FL-ita]:
Theorem .1.
(Theorem 1 in [sec24-defending-FL-ita]). For any random variable $D\in\mathbb{R}^{d}$ and $W\in\mathbb{R}^{m}$, we have
@@
By simply applying Theorem .1 in the privacy quantification in smashed data about raw input in split inference settings, we reach Theorem IV.2. For better understanding, we follow [sec24-defending-FL-ita] to give the proof details of Theorem IV.2 as follows:
@@
Lemma .2.
(Lemma 2 in [sec24-defending-FL-ita]). For any $d_{x}$-dimensional semi-positive definite matrix AA ($A\in\mathbb{R}^{d_{x}\times d_{x}}$), we have:
@@
Let $X$ be the raw input, $\hat{X}$ be the reconstructed input, and $x\sim p(X)$ we have:
@@
$=$ $\mathbb{E}_{x}H(\hat{X}|X=x)$ (19)
$\leq$ $\mathbb{E}_{x}\left[\frac{d_{x}}{2}\log(2\pi e)+\frac{1}{2}\log\det(Cov(\hat{X}|X=x))\right]$ (20)
$=$ $\frac{d_{x}}{2}\log(2\pi e)+\mathbb{E}_{x}\left[\frac{1}{2}\log\det(Cov(\hat{X}|X=x))\right]$ (21)
$\leq$ $\frac{d_{x}}{2}\log(2\pi e)+\mathbb{E}_{x}\left[\frac{d_{x}}{2}\log\frac{tr(Cov(\hat{X}|X=x))}{d_{x}}\right]$ (22)
[…]
The first inequality holds because the Gaussian distribution is the maximum entropy distribution among distributions with the same mean value and covariance. The second inequality depends on Lemma .2. The Equation 23 holds because,
@@
Lemma .3.
(Theorem 9 in [scarlett2019introductory]). Under the minimax estimation setup in Section 4.2, fix $\epsilon>0$, and index $V\in\mathcal{V}$ is drawn from a prior distribution $P_{V}$. $X_{V}=\{x_{v_{1}},\dots,x_{v_{|\mathcal{V}|}}\}$ drawn from a probability distribution $P_{X|V}$ be a finite subset of $\mathcal{X}$ such that
@@
In Lemma .3, the $X_{V}$ is identified by $V$, so we let both $V$ and $X_{V}$ be the raw input $X$ in split inference. In this case, given the $V=X$, we have $X_{V}=V=X$, where the $X_{V}$ is uniquely identified by $V$. Then, we let $Z$ be the smashed data $Z$ and both $\hat{X}$ and $\hat{V}$ be the reconstructed input $\hat{X}$ in split inference system, and we have the Markov chain
@@
Follow [noorbakhsh2024inf2guard], we assume $X$ is uniform over $\mathcal{X}$. Then by applying Lemma .3 and using $I(X;\hat{X})=H(\hat{X})-H(\hat{X}|X)$, we reach out Theorem 4.4. ∎
[Uncaptioned image] Ruijun Deng received her B.S. degree in Computer Science from Fudan University in 2022. Currently, she is pursuing the Master of Computer Science degree with Fudan University, China. Her research interests include federated learning, split learning, and AI security.
[Uncaptioned image] Zhihui Lu (Member, IEEE) is a Professor at the School of Computer Science, Fudan University. He received a Ph.D. computer science degree from Fudan University in 2004 and is a member of the IEEE and China Computer Federation’s service computing specialized committee. His research interests are cloud computing and service computing technology, big data architecture, edge computing, and IoT distributed systems. He has (co-)authored two books and more than 100 journal articles and conference papers in these areas.
[Uncaptioned image] Qiang Duan (Senior Member, IEEE) is currently a Professor with the College of Information Sciences and Technology, The Pennsylvania State University, Abington, PA, USA. His current research interests include network virtualization and softwarization, cognitive and autonomous networking, and edge computing-based ubiquitous intelligence. He has published three books and more than 100 research papers in these areas. He has served as an editor/associate editor for multiple research journals.
[Uncaptioned image] Shijing Hu received his B.S. in Computer Science from Peking University in 2021. Currently, he is pursuing doctor’s degree at the School of Computer Science, Fudan University, China. His research interests include edge intelligence and large language models.